A horizontal force of 0.80N is required to pull a 5kg block across a table top at a constant speed. With the block initially at rest, a 20g bullet fired horizontally into the block causes the block to slide 1.5m before coming to rest again. Determine the speed v of the bullet , where the bullet is assumed to be embedded in the block.
1
Expert's answer
2021-08-23T12:40:43-0400

Given:

"M=5.0\:\rm kg"

"m=0.02\:\rm kg"

"F=0.80\:\rm N"

"d=1.5\:\rm m"


The law of conservation of momentum gives

"mu=(M+m)v"

The energy-work theorem gives

"\frac{(M+m)v^2}{2}=Fd"

Finally, the initial speed of a bullet

"u=\frac{M+m}{m}\sqrt{\frac{2Fd}{M+m}}"

"u=\frac{5.0+0.02}{0.02}\sqrt{\frac{2*0.8*1.5}{5.0+0.02}}=174\:\rm m/s"